OpenStudy (anonymous):

Multiply or divide.5x+10/x+2•2x/4x-10

OpenStudy (jack1):

both

OpenStudy (jack1):

also, add some brackets so the q's a bit clearer ie is it: (5x+10)/(x+2)•(2x)/(4x-10) or 5x+(10/x+2)•(2x/4x) -10

OpenStudy (jack1):

is makes a difference to the final answer

OpenStudy (jack1):

sorry, different problem, back

OpenStudy (jack1):

so: (5x+10)/(x+2) times (2x/(4x-10) yeah?

OpenStudy (jack1):

so when multiplying fractions, times numerator by numerator and denominator by denominator so top x top = (5x+10) times 2x = (5x +10)(2x)

OpenStudy (jack1):

botom times bottom = (x+2) times (4x-10) =(x +2)(4x-10)

OpenStudy (jack1):

so your long answer is : (5x +10)(2x) / (x +2)(4x-10) but you can further simplify this:

OpenStudy (jack1):

(5x +10)(2x) becomes 10x^2 +20x (x+2)(4x+10) becomes 4x^2 + 18x +20

OpenStudy (jack1):

now, you can take 2 out of top and bottom, so it becomes: 2(5x^2 +10x) / 2(2x^2 + 9x +10)

OpenStudy (jack1):

= (5x^2 +10x) /(2x^2 + 9x +10)

OpenStudy (jack1):

thats about it

OpenStudy (anonymous):

That's the final answer right?

OpenStudy (jack1):

yeah i cant see a way to reduce any further

OpenStudy (anonymous):

Thnx for ur help I truly appreciated

OpenStudy (jack1):

welcome, no worries